# A Second Main Theorem for Entire Curves Intersecting Three Conics

### v1

@@ 6.2. Geometric Interpretation of Wronskians
Let $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ be a normal crossing divisor of three smooth conics in $\mathbb{P}^{2}$, defined by homogeneous quadratic polynomials $A(\mathbf{Z})$, $B(\mathbf{Z})$, and $C(\mathbf{Z})$, where $\mathbf{Z}=(Z_{0},Z_{1},Z_{2})$.
On the affine chart $U_{0}\coloneqq\{Z_{0}\neq 0\}$ with coordinates $(x_{1},x_{2})=(Z_{1}/Z_{0},Z_{2}/Z_{0})$, let $a(x_{1},x_{2})$, $b(x_{1},x_{2})$, and $c(x_{1},x_{2})$ be the dehomogenized polynomials of $A,B,C$. Denote $a_{i}=\partial a/\partial x_{i}$ and $b_{j}=\partial b/\partial x_{j}$ for $i,j=1,2$.
@@
On $U_{0,C}$, noting that $c$ is nowhere vanishing, the logarithmic $1$-forms $(\mathrm{log}\,a)^{\prime}$ and $(\mathrm{log}\,b)^{\prime}$ generate the logarithmic cotangent bundle $T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$. Let $f\colon(\mathbb{C},0)\ni t\mapsto f(t)\in U_{0,C}\subset\mathbb{P}^{2}$ be a germ of a holomorphic curve, with local representation $f=(f_{1},f_{2})$ in the coordinates $(x_{1},x_{2})$. Denote $a(f(t))\coloneqq(f^{*}a)(t)=a(f_{1}(t),f_{2}(t))$ and $b(f(t))\coloneqq(f^{*}b)(t)=b(f_{1}(t),f_{2}(t))$.
@@ 6.3. A Refined Estimate
The following estimate, formulated in the notation of Subsection 6.1, is a key step in proving our main theorem.
@@
Our proof builds on the geometric interpretation of Wronskian operators and incorporates techniques from the proof of [Huynh-Vu-Xie-2017, Theorem 3.1]. This approach allows us to improve the constant on the right-hand side from $\frac{3\tilde{m}}{\tilde{t}}$ (check Theorem 1.2) to $\frac{2\tilde{m}}{\tilde{t}}$. This refinement constitutes a key contribution of the present work and underscores the advantage of using invariant jet differentials over normal ones.
@@
Note that for $k=1,2$, the pole order of $\left(\mathrm{log}\,a(g)\right)^{(k)}$ (resp. $\left(\mathrm{log}\,b(g)\right)^{(k)}$) at $z\in g^{-1}(U_{0,C})\subset\mathbb{C}$ is at most $k\min\{\operatorname{ord}_{z}A(g),1\}$ (resp. $k\min\{\operatorname{ord}_{z}B(g),1\}$). Thus
$D_{\sigma,g_{[2]}}|_{g^{-1}(U_{0,C})}\leqslant\tilde{m}\Big(\sum_{z\in g^{-1}(U_{0,C})}2\big(\min\{\operatorname{ord}_{z}A(g),1\}+\min\{\operatorname{ord}_{z}B(g),1\}\big)\,z\Big).$
@@
$D_{\sigma,g_{[2]}}\leqslant 2\tilde{m}\sum_{z\in\mathbb{C}}\Big(\sum_{Q\in\{A,B,C\}}\min\{\operatorname{ord}_{z}Q(g),1\}\Big)z.$
@@ 6.4. An Effective Constant $c$ in the Second Main Theorem for Three Conics
satisfying $0<t<m$ and such that the divisor $Z$ of $\omega_{1}$ on the second level $X_{2}$ of the logarithmic Demailly–Semple tower for $(\mathbb{P}^{2},\mathcal{C})$ is irreducible and reduced modulo $\Gamma_{2}$. Then, by the argument in Section 6.1, for any rational number $\tau\in(0,\tau_{1}(t/m))$, we obtain a nontrivial global section
@@
Meanwhile, in Case 2, to ensure that $\omega_{2}\in H^{0}\big(\mathbb{P}^{2},\,E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(7-t)\big)$ satisfies $m/(t-7)\leqslant c$, the positive integers $m$ and $t$ associated to the first family $\omega_{1}$ must satisfy $t\geqslant\frac{m}{c}+7.$
Taking $c=19$, we obtain the desired estimate (1.5) for positive integer pairs $(m,t)$ satisfying $0<t<\frac{1940}{2109}m$ or $\quad t\geqslant\frac{m}{19}+7.$ Note that $\frac{1940}{2109}\times 8\approx 7.36$ and $\frac{1940}{2109}\times 7<7$. To exclude the exceptional cases of $(m,t)$, it suffices to prove the Key Vanishing Lemma I.
@@
7. Proof of the Key Vanishing Lemmas I and II
@@ 7.1. Basic Calculations
Under generic coefficient assumptions, the Jacobian curve $\{J=0\}$ satisfies the following transversality properties (cf. e.g. [Guo-Sun-Wang-2022, CHSX2025]):
@@
    Transverse intersection with each defining conic $\mathcal{C}_{i}$ for $i=1,2,3$
@@
    Transverse intersection with each coordinate hyperplane $\{Z_{i}=0\}$ for $i=0,1,2$
@@
Let $\mathcal{H}:=\{Z_{0}=0\}+\{Z_{1}=0\}+\{Z_{2}=0\}$ be a divisor consisting of the three coordinate hyperplanes in $\mathbb{P}^{2}$. Then the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{H})$ is globally generated by the sections $(\mathrm{log}\,\frac{Z_{0}}{Z_{2}})^{\prime}$ and $(\mathrm{log}\,\frac{Z_{1}}{Z_{2}})^{\prime}$.
Proof.
The verification is straightforward by direct computation on the three affine charts $\{Z_{i}\neq 0\}$, $i=0,1,2$. ∎
@@
Proof.
By a similar direct computation on each affine chart $\{Z_{i}\neq 0\}$ as in Proposition 7.1. ∎
@@
Thus sections $\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}$ and $\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}$ continue to generate the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ in a neighborhood of $\{J\neq 0\}\cap\{C=0\}$ inside $\{J\neq 0\}$, as the transformation matrix remains regular when approaching this locus. The analogous argument holds on charts $U_{1}=\{Z_{1}\neq 0\}$ and $U_{2}=\{Z_{2}\neq 0\}$, completing the proof. ∎
@@
Our carefully chosen basis (7.5) induces a highly homogeneous representation of the logarithmic jet differentials, which substantially mitigates the complications arising from inhomogeneous coordinate transitions on affine charts.
@@ 7.2. Exemplary Case $m=3,t=1$
    Employing algebro-geometric techniques to establish the uniform bound of $\ell$, thereby reducing the infinite-dimensional problem to a computationally tractable finite-dimensional linear algebra framework;
@@
    Implementing a verification scheme (executed in Maple) that confirms the identical vanishing $P\equiv 0$ through explicit computation with judiciously chosen specializations of the coefficients $A_{\bullet}$, $B_{\bullet}$, and $C_{\bullet}$.
The semi-continuity principle then guarantees the extension of this vanishing property to all generic coefficient choices $A_{\bullet}$, $B_{\bullet}$, $C_{\bullet}$.
@@
(7.7) $P=R_{0}(x_{1},x_{2})\begin{vmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{a}{c}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime\prime}\end{vmatrix}+\sum_{k=0}^{3}R_{0}^{k,3-k}(x_{1},x_{2})\left(\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}\right)^{k}\left(\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\right)^{3-k},$
@@
Claim. The exponent satisfies $\ell\leqslant 1$.
@@
  • The vector $\Psi^{-1}(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}=\frac{1}{D^{2}}(\cdots)^{\mathsf{T}}$ introduces a double pole along $\{D=0\}$ by direct computations. The key aspect of our argument is to bound the pole orders along $\{D=0\}$ by effective small numbers (not necessary optimal), thereby managing the computational complexity through computer-assisted calculations subsequently.
@@
(7.18) $r_{0}=x_{2}\left(\sum_{\begin{subarray}{c}p+q\leqslant 2\end{subarray}}\lambda_{p,q}x_{1}^{p}x_{2}^{q}\right),\quad r_{0}^{k,3-k}=x_{2}\bigg(\sum_{\begin{subarray}{c}p+q\leqslant 8\end{subarray}}\lambda_{p,q}^{k,3-k}x_{1}^{p}x_{2}^{q}\bigg)\quad(\lambda_{p,q},\lambda_{p,q}^{k,3-k}\in\mathbb{C}).$
@@
Indeed, we can confirm this conclusion through Maple computations with specific three conics:

### v2

@@ 6.2. Geometric Interpretation of Wronskians
Let $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ be a normal crossing divisor of three smooth conics in $\mathbb{P}^{2}$, defined by homogeneous quadratic polynomials $A(\mathbf{Z})$, $B(\mathbf{Z})$, and $C(\mathbf{Z})$ in $\mathbb{C}[Z_{0},Z_{1},Z_{2}]$, where $\mathbf{Z}=(Z_{0},Z_{1},Z_{2})$.
On the affine chart $U_{0}\coloneqq\{Z_{0}\neq 0\}\subset\mathbb{P}^{2}$ with coordinates $(x_{1},x_{2})=(Z_{1}/Z_{0},Z_{2}/Z_{0})$, let $a(x_{1},x_{2})$, $b(x_{1},x_{2})$, and $c(x_{1},x_{2})$ be the dehomogenized polynomials of $A,B,C$ (see Subsection 7.1). Denote $a_{i}=\partial a/\partial x_{i}$ and $b_{j}=\partial b/\partial x_{j}$ for $i,j=1,2$.
@@
On $U_{0,C}$, noting that $c$ is nowhere vanishing, the logarithmic $1$-forms $(\mathrm{log}\,a)^{\prime}$ and $(\mathrm{log}\,b)^{\prime}$ generate the logarithmic cotangent bundle $T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$. Let $f\colon(\mathbb{C},0)\ni t\mapsto f(t)\in U_{0,C}\subset\mathbb{P}^{2}$ be a germ of a holomorphic curve, with local representation $f=(f_{1},f_{2})$ in the coordinates $(x_{1},x_{2})$. Denote
(6.7) $a(f(t))\coloneqq a(f_{1}(t),f_{2}(t))\;\text{ and }\;b(f(t))\coloneqq b(f_{1}(t),f_{2}(t)).$
@@ 6.3. A Refined Estimate
Here is a key step in proving our main theorem.
@@
Our proof builds on the geometric interpretation of Wronskian operators and incorporates techniques from the proof of [Huynh-Vu-Xie-2017, Theorem 3.1]. This approach allows us to improve the constant on the right-hand side from $\frac{3\tilde{m}}{\tilde{t}}$ (cf. Theorem 1.2) to $\frac{2\tilde{m}}{\tilde{t}}$. This refinement constitutes a key contribution of the present work and underscores the advantage of using invariant logarithmic $2$-jet differentials over the Green-Griffiths ones.
@@
Note that for $k=1,2$, the pole order of $\left(\mathrm{log}\,a(g)\right)^{(k)}$ (resp. $\left(\mathrm{log}\,b(g)\right)^{(k)}$) at $z\in g^{-1}(U_{0,C})\subset\mathbb{C}$ is at most $k\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{1},1\}$ (resp. $k\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{2},1\}$). Thus
$D_{\sigma,g_{[2]}}|_{g^{-1}(U_{0,C})}\leqslant\tilde{m}\Big(\sum_{z\in g^{-1}(U_{0,C})}2\big(\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{1},1\}+\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{2},1\}\big)\,z\Big).$
@@
$D_{\sigma,g_{[2]}}\leqslant 2\tilde{m}\sum_{z\in\mathbb{C}}\Big(\sum_{i=1}^{3}\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{i},1\}\Big)z.$
@@ 6.4. An Effective Constant $c$ in the Second Main Theorem for Three Conics
satisfying $0<t<m$ and such that the divisor $Z$ of $\omega_{1}$ on the second level $X_{2}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$ is irreducible and reduced modulo $\Gamma_{2}$. Then, by the argument in Section 6.1, for any rational number $\tau\in(0,\tau_{1}(t/m))$, we obtain a nontrivial global section
@@
Meanwhile, in Case 1, to ensure that $\omega_{2}\in H^{0}\big(\mathbb{P}^{2},\,E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(7-t)\big)$ satisfies $m/(t-7)\leqslant c$, the positive integers $m$ and $t$ associated to the first family $\omega_{1}$ must satisfy $t\geqslant\frac{m}{c}+7.$
Taking $c=19$, we obtain the desired estimate (1.5) for positive integer pairs $(m,t)$ satisfying $0<t<\frac{1940}{2109}m$ or $t\geqslant\frac{m}{19}+7.$ Note that $\frac{1940}{2109}\times 8\approx 7.36$ and $\frac{1940}{2109}\times 7<7$. To exclude the exceptional cases of $(m,t)$, it suffices to prove the Key Vanishing Lemma I.
@@
7. Proof of the Key Vanishing Lemmas
@@ 7.1. Basic Calculations
Under generic coefficient assumptions, the Jacobian curve $\{J=0\}$ satisfies the following transversality properties (cf., e.g., [Guo-Sun-Wang-2022, CHSX2025]):
@@
    Transverse intersection with each conic $\mathcal{C}_{i}$ for $i=1,2,3$.
@@
    Transverse intersection with each coordinate hyperplane $\{Z_{i}=0\}$ for $i=0,1,2$.
@@
We begin with the following two results, which follow from direct computation on the affine charts $\{Z_{i}\neq 0\}\subset\mathbb{P}^{2}$, $i=0,1,2$.
@@
Let $\mathcal{H}:=\{Z_{0}=0\}+\{Z_{1}=0\}+\{Z_{2}=0\}$ be a divisor consisting of the three coordinate hyperplanes in $\mathbb{P}^{2}$. Then the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{H})$ is globally generated by the sections $(\mathrm{log}\,\frac{Z_{0}}{Z_{2}})^{\prime}$ and $(\mathrm{log}\,\frac{Z_{1}}{Z_{2}})^{\prime}$. ∎
@@
Thus the sections $\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}$ and $\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}$ continue to generate the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ in a neighborhood of $\{J\neq 0\}\cap\{C=0\}$ inside $\{J\neq 0\}$, as the transformation matrix remains regular when approaching this locus. The analogous argument holds on charts $U_{1}=\{Z_{1}\neq 0\}$ and $U_{2}=\{Z_{2}\neq 0\}$, completing the proof. ∎
@@
Our carefully chosen basis (7.5) induces a highly homogeneous representation of the logarithmic jet differentials, which substantially mitigates the complications arising from inhomogeneous coordinate transitions on affine charts.333 We also attempted two other natural frame systems on $\{Z_{0}A\neq 0\}$ and $\{Z_{1}B\neq 0\}$, equipped with the obvious local frames, and applied the same strategy. Unfortunately, the computational complexity “explodes” beyond the capacity of current standard computers.
@@ 7.2. Exemplary Case $m=3,t=1$
    Employing algebro-geometric techniques to establish a uniform upper bound of $\ell$, thereby reducing the infinite-dimensional problem to a computationally tractable finite-dimensional linear algebra framework;
@@
    Implementing a verification scheme (executed in Maple) that confirms the identical vanishing $P\equiv 0$ through explicit computation with certain carefully chosen coefficients $A_{\bullet}$, $B_{\bullet}$, and $C_{\bullet}$.
The semi-continuity argument then guarantees the extension of this vanishing property to all generic coefficients $A_{\bullet}$, $B_{\bullet}$, $C_{\bullet}$.
@@
(7.7) $P=R_{0}(x_{1},x_{2})\begin{vmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{a}{c}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime\prime}\end{vmatrix}+\sum_{k=0}^{3}R_{0}^{k,3-k}(x_{1},x_{2})\left(\Big(\mathrm{log}\,\frac{a}{c}\Big)^{\prime}\right)^{k}\left(\Big(\mathrm{log}\,\frac{b}{c}\Big)^{\prime}\right)^{3-k},$
@@
Key Observation 7.13.
The exponent satisfies $\ell\leqslant 1$.
@@
  • The vector $\Psi^{-1}(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}=\frac{1}{D^{2}}(\cdots)^{\mathsf{T}}$ produces a double pole along $\{D=0\}$ as a direct (Maple) computation shows. The essential step in our argument is to bound the pole orders along $\{D=0\}$ by explicit (though not necessarily sharp) effective exponents. This translates the problem into a concrete computational one. Since the required computations remain within the reach of contemporary machines, it becomes solvable. The existence of such an exponent bound would be foreshadowed and predicted by the well-known finiteness of cohomology $\dim H^{0}(\cdots)<\infty$.
@@
(7.19) $r_{0}=x_{2}\sum_{\begin{subarray}{c}p+q\leqslant 2\end{subarray}}\lambda_{p,q}x_{1}^{p}x_{2}^{q},\quad r_{0}^{k,3-k}=x_{2}\sum_{\begin{subarray}{c}p+q\leqslant 8\end{subarray}}\lambda_{p,q}^{k,3-k}x_{1}^{p}x_{2}^{q}\qquad(\lambda_{p,q},\lambda_{p,q}^{k,3-k}\in\mathbb{C}).$
@@
Indeed, this conclusion can be confirmed through explicit computer-assisted computations for the following specific three conics: